OpenStudy (anonymous):

A 20-foot ladder is leaning against the side of a building. The bottom of the ladder is 4 feet from the wall. How many feet above the ground does the ladder touch the wall? ANSWER CHOICES A. 16 FT B. √384 FT C. 12 FT D. √ 416 FT

OpenStudy (ghazi):

\[(20)^2=4^2+x^2\]

OpenStudy (anonymous):

Thanks.

OpenStudy (ghazi):

:D YW

OpenStudy (anonymous):

let me try to solve this.

OpenStudy (ghazi):

sure :)

OpenStudy (anonymous):

Hmm.

OpenStudy (ghazi):

what?

OpenStudy (anonymous):

let me give u the available answer choices

OpenStudy (anonymous):

I think its B

OpenStudy (ghazi):

400=x^2+16 400-16=x^2 x=19.59